# MARS: Unleashing the Power of Variance Reduction for Training Large Models

### v1

@@ 4.1 Experimental Setup
(a) Training Loss
(b) Validation Loss
(c) Wall-clock time
Figure 4: The accuracy for 0-shot evalution on Hellaswag (Zellers et al., 2019) dataset for AdamW, Muon and MARS on GPT-2 medium (355M) and large (770M) models.
4.2 Results
In Figures 1, 2, and 3, we demonstrate the training and validation losses as a function of training tokens and wall-clock time for various model sizes [^1] The training loss curves are smoothed using an Exponential Moving Average.. Across the small, medium, and large GPT-2 models, MARS consistently surpasses both the AdamW and Muon baselines in training and validation losses. The performance gap becomes more pronounced with increasing model size. Notably, MARS exhibits both rapid initial decay and sustained superiority throughout the training process. Further, we explore the performance of additional learning rate choices in Appendix B. Notably, the best validation loss of MARS achieved in our GPT-2 small experiments is 2.852. For comparison, Sophia-H (Liu et al., 2023) achieves a validation loss of 2.901, and Sophia-G (Liu et al., 2023) achieves 2.875 under the same settings and datasets as ours. These results demonstrate that our reported performance is highly competitive with state-of-the-art optimizers.
In Figure 1(c), 2(c), and 3(c), we compare the wall-clock time of different algorithms. We observe that our algorithm, MARS, has a slightly higher per-iteration cost compared to AdamW but is much faster than Muon. Additionally, MARS consistently demonstrates lower validation losses than both AdamW and Muon within equivalent training durations.
We also test our optimization algorithm on the downstream task of commonsense natural language inference using the Hellaswag dataset (Zellers et al., 2019), evaluated with 00-shot. The results, shown in Figure 4, are obtained from GPT-2 medium and large models pre-trained for 100,000 steps (50B tokens). The models pre-trained with MARS outperformed those pre-trained with AdamW and Muon optimizers, validating an enhanced downstream performance within the same number of pre-training steps.
4.3 MARS and MARS-approx.
@@ 4.2 Results
We then conduct experiments to compare the performance of MARS and MARS-approx (MARS-AdamW version) on GPT-2 small and medium models, the results are shown in Figure 5. Models trained with MARS exhibit slightly better performance than those trained with MARS-approx, though the performance gap in terms of training tokens is marginal. This suggests that the approximate version is a practical alternative in scenarios where computational cost is a priority, as it introduces only minimal performance loss. However, in settings where achieving higher validation accuracy is critical, the exact version is recommended.
@@
Figure 5: Validation loss curves for MARS and MARS-approx on GPT-2 small (125M, left) and medium (355M, right).
5 Related Work
Adaptive Gradient Methods.
RProp (Riedmiller and Braun, 1993) is probably one of the earliest adaptive gradient methods by dynamically adjusting the learning rate. AdaGrad (Duchi et al., 2011; McMahan and Streeter, 2010) adjusts the learning rate based on the geometry of the training data observed during earlier iterations. To tackle with the issue of diminishing gradient in AdaGrad, Tieleman (2012) introduced RMSProp by incorporating the idea of exponential moving average. A significant advancement came with Adam (Kingma, 2014), which integrated RMSProp with Nesterov’s momentum (Nesterov, 1983, 2013) achieving superior performance and becoming a prevalent optimizer in deep neural network training. Later, Loshchilov (2017) proposed to decouple weight decay from gradient calculations in Adam and introduced AdamW, an optimization algorithm having become the predominant optimization algorithm in contemporary deep learning applications. To fix the convergence issue of Adam, Reddi et al. (2019) introduced the AMSGrad optimizer, which maintains a running maximum of past second-order momentum terms to achieve non-increasing step sizes. Subsequently, Chen et al. (2018) unified AMSGrad and SGD within the Padam framework by introducing a partial adaptive parameter to control the degree of adaptiveness. Notably, AdamW and its variations have been widely used in the training of popular large language models, including OPT (Zhang et al., 2022), Llama 3 (Dubey et al., 2024), and DeepSeek-V2 (Liu et al., 2024).
Variance Reduction Methods.
Variance Reduction for Adaptive Gradient Methods.
Few works have explored the application of variance reduction techniques to adaptive gradient methods. To the best of our knowledge, the only exceptions are Adam+ and SuperAdam. Adam+ (Liu et al., 2020) attempts to reduce the variance of first-order moment in Adam by estimating the gradient only at extrapolated points. SuperAdam (Huang et al., 2021), an adaptive gradient algorithm that integrates variance reduction with AdamW to achieve improved convergence rates. However, these variance-reduced adaptive gradient methods have primarily been validated on basic computer vision tasks, such as MNIST (Schölkopf and Smola, 2002) and CIFAR-10 (Krizhevsky et al., 2009), and simple natural language modeling tasks, like SWB-300 (Saon et al., 2017), using straightforward architectures such as LeNet (LeCun et al., 1998), ResNet-32 (He et al., 2016), 2-layer LSTMs (Graves and Graves, 2012), and 2-layer Transformers (Vaswani, 2017). As a result, a significant gap remains in the successful application of variance reduction techniques to adaptive gradient methods, particularly in the rapidly evolving domain of large language models.
6 Conclusion
Acknowledgments
We thank Yuan Cao, Yiming Dong, and Zhuoqing Song for the valuable discussions throughout various stages of this work.
Appendix A Proofs of Lemma 3.4
Proof of Lemma 3.4.
Substituting (3.23) into (3.24), we obtain
$\mathbf{m}_{t}=b_{1}\big{(}a_{1}\mathbf{u}_{t-1}+a_{2}\mathbf{g}_{t}\big{)}+b_{2}\mathbf{g}_{t}=a_{1}b_{1}\mathbf{u}_{t-1}+(b_{1}a_{2}+b_{2})\mathbf{g}_{t}.$ (A.1)
On the other hand, shifting the index of (3.24) by $1$, we have
$\mathbf{m}_{t-1}=b_{1}\mathbf{u}_{t-1}+b_{2}\mathbf{g}_{t-1}.$ (A.2)
Combining (A.1) and (A.2), we obtain the iterative update of $\mathbf{m}_{t}$ from its previous value $\mathbf{m}_{t-1}$ as:
$\mathbf{m}_{t}$ $=a_{1}\mathbf{m}_{t-1}-a_{1}b_{2}\mathbf{g}_{t-1}+(b_{1}a_{2}+b_{2})\mathbf{g}_{t}$
$=a_{1}\mathbf{m}_{t-1}+(b_{1}a_{2}-a_{1}b_{2}+b_{2})\mathbf{g}_{t}+a_{1}b_{2}(\mathbf{g}_{t}-\mathbf{g}_{t-1}).$
This completes the proof. ∎
Appendix B Additional Experiments
B.1 Learning Rates
In our experiments, we explore various learning rate choices on GPT-2 small. AdamW with a learning rate of $6\times 10^{-4}$ demonstrates superior performance in terms of validation loss during the initial 37 billion tokens of training. However, a higher learning rate of $3\times 10^{-3}$ yields significantly better results in the final 13 billion tokens. We then further extend our experiments by comparing AdamW and MARS-AdamW under identical hyperparameters, both using the higher learning rate of $3\times 10^{-3}$. The results, presented in Figure 6, show that even against an alternative baseline with hyperparameters optimized for AdamW, MARS-AdamW consistently outperforms AdamW. We observe that there is still potential for further parameter tuning and performance improvements with MARS.
Figure 6: Validation loss with respect to training tokens for AdamW with learning rates $6\times 10^{-4}$, $3\times 10^{-3}$ and MARS with learning rate $3\times 10^{-3}$ on GPT-2 small model (125M). The model trained with MARS achieves a validation loss of 2.852.
B.2 Comparison between MARS and Lion
The key difference between MARS-AdamW and MARS-Lion lies in their precondition matrix. We evaluate MARS-AdamW against MARS-Lion under identical hyperparameters on GPT-2 small. As shown in Figure 7, MARS-AdamW outperforms MARS-Lion in both training and validation loss. Therefore, we mainly use MARS-AdamW in the rest of our experiments.
(a) Training Loss
(b) Validation Loss
@@ B.2 Comparison between MARS and Lion
Figure 7: The training and validation loss curves, plotted against training tokens for Lion and MARS on GPT-2 small model (125M).
B.3 Sensitivity to $\gamma$.
To explore the impact of $\gamma_{t}$, we tested various $\gamma_{t}$ schedules, including constant and linearly changing schedules. Other hyper-parameters are set the same as the main experiments, and the results are summarized in Table 1. Specifically, the constant $\gamma$ values of 0.025 and 0.05 exhibit the best performance among the $\gamma_{t}$-schedules evaluated. We are using a $\gamma_{t}$ schedule of $0.025$ across all our other experiments.
Table 1: The validation loss for GPT-2 small (125M) model with different $\gamma$-schedules.
$\gamma$-schedule $\gamma$ Validation loss
Constant 0.025 2.871
0.05 2.871
Changing 0.025$\rightarrow$0.05 2.872
0.1$\rightarrow$0 2.873
0$\rightarrow$0.1 2.891
0$\rightarrow$0.2 2.899
Appendix C Hyper-parameter Choices
Table 2 summarizes the architectural hyperparameters for GPT-2 models with 125M (small), 355M (medium), and 770M (large) parameters. Table 3 lists the general hyperparameters used across all experiments, while Tables 4, 5 and 6 present the training hyperparameters for the small, medium, and large models, respectively.
@@ Appendix C Hyper-parameter Choices
Table 4: Hyper-parameters for GPT-2 small experiment. We use $\gamma_{t}\equiv 0.025$ for MARS.
Hyper-parameter AdamW Muon MARS
Max learning rate 6e-4 2e-2 6e-3
Min learning rate 3e-5 3e-5 3e-5
$(\beta_{1},\beta_{2})$ (0.9, 0.95) $\mu=0.95$ (0.95, 0.99)
Table 5: Hyper-parameters for GPT-2 medium experiment. We use $\gamma_{t}\equiv 0.025$ for MARS.
Hyper-parameter AdamW Muon MARS
Max learning rate 3e-4 1e-2 3e-3
Min learning rate 6e-5 6e-5 6e-5
$(\beta_{1},\beta_{2})$ (0.9, 0.95) $\mu=0.95$ (0.95, 0.99)
Table 6: Hyper-parameters for GPT-2 large experiment. We use $\gamma_{t}\equiv 0.025$ for MARS.
Hyper-parameter AdamW Muon MARS
Max learning rate 2e-4 6.67e-3 2e-3
Min learning rate 1e-5 1e-5 1e-5
$(\beta_{1},\beta_{2})$ (0.9, 0.95) $\mu=0.95$ (0.95, 0.99)

### v2

@@ D.1 Supplementary Results for the Main Experiments
D.2 MARS and MARS-approx
@@
We then conduct experiments to compare the performance of MARS and MARS-approx (MARS-AdamW instantiation) on GPT-2 small and medium models, the training and validation loss curves are shown in Figures 8 and 9. Models trained with MARS exhibit consistently better performance than those trained with MARS-approx. This suggests that: (a) The exact version, which employs the variance reduction formulation, is more fundamental than the approximate version. (b) The approximate version serves as a practical alternative in scenarios where computational efficiency is a priority, as it incurs only minimal performance loss. However, in settings where maximizing validation accuracy is crucial, the exact version is recommended.
@@
Figure 8: Training loss curves for MARS-AdamW and MARS-AdamW-approx on GPT-2 small (125M, left) and medium (355M, right), pretrained with OpenWebText dataset and plotted against training tokens.
Figure 9: Validation loss curves for MARS-AdamW and MARS-AdamW-approx on GPT-2 small (125M, left) and medium (355M, right), pretrained with OpenWebText dataset.
D.3 Experiments on FineWeb-Edu 100B Dataset
FineWeb-Edu dataset (Lozhkov et al., 2024) is a high-quality dataset based on well-filtered educational web pages. To better investigate the efficiency of our algorithm, we also use FineWeb-Edu 100B, a subset of FineWeb-Edu with around 100B tokens to train GPT-2 small (125M) and XL (1.5B, with the same learning rates as GPT-2 large models) models with optimizers including AdamW, Muon and MARS-AdamW-approx. We leave around 0.1B tokens for validation and other tokens for training. The training and evaluation curves are shown in Figures 10 and 11. It can be seen that our algorithms can also achieve better performances even with different datasets. For a comprehensive investigation, we evaluate these models on metrics same as experiences in Section 5.2, and the results are shown in Tables 7 and 8. We also compare the results with the open-source GPT-2 models on Hugging Face Radford et al. (2019) (denoted as “OpenAI-Comm.” in the tables). Compared with Table 2, it can be observed that this dataset is actually better for the superior performances. However, models trained with our algorithm can also show advantages over baseline optimization approaches trained with such a high-quality dataset.
Table 7: The evaluation results of small models pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSwag = HellaSwag, WG = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.
OpenAI-Comm. 39.48 22.70 48.72 31.14 27.20 62.51 51.62 22.92 64.40 41.19
AdamW 51.43 26.54 55.78 36.26 30.60 64.53 50.36 24.49 71.50 45.72
Muon 47.85 27.56 57.16 33.46 31.60 63.66 51.30 23.17 67.30 44.78
MARS-AdamW 52.23 27.39 55.84 36.91 32.20 64.80 49.96 22.95 71.10 45.93
Table 8: The evaluation results of XL models pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSwag = HellaSwag, WG = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.
OpenAI-Comm. 51.05 28.50 61.77 50.89 32.00 70.51 58.33 25.24 76.00 50.48
AdamW 68.22 38.40 61.13 53.93 39.00 72.69 54.78 25.47 85.30 55.43
Muon 64.18 36.52 58.38 51.83 37.40 72.03 55.56 24.93 81.90 53.64
MARS-AdamW 66.54 39.85 63.82 56.52 41.20 73.34 56.59 23.86 86.00 56.41
@@ D.3 Experiments on FineWeb-Edu 100B Dataset
Figure 10: Training loss curves for AdamW, Muon and MARS-AdamW-approx on GPT-2 small (125M, left) and XL (1.5B, right), pretrained with FineWeb-edu 100B dataset and plotted against training tokens.
Figure 11: Validation loss curves for AdamW, Muon and MARS-AdamW-approx on GPT-2 small (125M, left) and XL (1.5B, right), pretrained with FineWeb-edu 100B dataset and plotted against training tokens.
D.4 Computer Vision Experiments
We also carry out experiments on the classification task in the field of computer vision. We conduct experiments with ResNet-18 model (He et al., 2016) on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009) for AdamW, Lion, Shampoo [^3] In practice, we use Distributed Shampoo (Shi et al., 2023) to facilitate the training of Shampoo., Muon, and variants of MARS instantiation, following the setting in Chen et al. (2018a).
We do grid research to explore the best hyper-parameters for each of these optimization methods. We search over $\{10^{-5},...,10^{0}\}$ for the learning rate and $\{0,...,1.0\}$ for the weight decay. We set $\beta_{1}=0.9$ and search over $\{0.99,0.999\}$ for $\beta_{2}$ for AdamW and Muon; fix $\beta_{1}=0.9$ and search $\beta_{2}$ over $\{0.99,0.999\}$ for Lion; search over $\{0.9,0.95\}$ for $\beta_{1}$ and $\{0.95,0.99,0.999\}$ for $\beta_{2}$ for Shampoo; and we fix the $(\beta_{1},\beta_{2})=(0.95,0.99)$ and $\gamma=0.025$ for MARS models. We train for 200 epochs with training batch size 128 on 1 NVIDIA A6000 GPU. And we also apply MultiStepLR scheduler so that the learning rate would decrease to 10% of the original rate at the 100th epoch and to 1% at the 150th epoch. We display the test loss and test accuracy for CIFAR-10 and CIFAR-100 datasets in Figures 12 and 13, respectively. The results show that our algorithm can achieve better validation loss after the decay of learning rate and better test accuracy within the final stage of training.
(a) Test Loss
(b) Test Accuracy
Figure 12: The test loss and test accuracy for different optimizers on CIFAR-10 dataset.
(a) Test Loss
(b) Test Accuracy
Figure 13: The test loss and test accuracy for different optimizers on CIFAR-100 dataset.
We also compare the performances among the baselines of AdamW, Lion and Shampoo without variance reduction, the approximate and the exact versions of MARS instantiations. The test loss and accuracy curves for CIFAR10 dataset are shown in Figures, and Figures, respectively. And the test loss and accuracy curves for CIFAR100 dataset are shown in Figures 1415, and Figures 1617, respectively. Moreover, we list the best test losses and accuracies in Table 9. It can be observed that the exact versions perform a little better than the approximate versions, but much better than the baseline approaches, showing the superiority of variance reduction in MARS.
Table 9: The best test losses and accuracies of different optimizers on CIFAR10 and CIFAR100 datasets during the training epochs. The bolded are the best test loss or best test accuracy among the listed optimizers.
Optimizer Best Test Loss Best Test Accuracy
CIFAR10 CIFAR100 CIFAR10 CIFAR100
AdamW 0.230 1.726 94.81 73.70
Lion 0.245 1.351 94.68 74.28
Shampoo 0.354 2.426 94.65 74.27
Muon 0.306 2.608 95.08 74.64
MARS-AdamW-approx 0.199 0.971 95.29 76.97
MARS-AdamW 0.193 0.888 95.26 77.38
MARS-Lion-approx 0.202 0.985 95.05 75.97
MARS-Lion 0.219 0.991 94.98 76.15
MARS-Shampoo-approx 0.202 1.256 94.92 74.80
MARS-Shampoo 0.194 0.982 94.98 75.83
(a) AdamW-series
(b) Lion-series
(c) Shampoo-series
Figure 14: The test loss curves for the baselines of AdamW, Lion and Shampoo without variance reduction, the approximate and the exact versions of MARS instantiations on CIFAR-10 dataset.
(a) AdamW-series
(b) Lion-series
(c) Shampoo-series
Figure 15: The test accuracy curves for the baselines of AdamW, Lion and Shampoo without variance reduction, the approximate and the exact versions of MARS instantiations on CIFAR-10 dataset.
(a) AdamW-series
(b) Lion-series
(c) Shampoo-series
Figure 16: The test loss curves for the baselines of AdamW, Lion and Shampoo without variance reduction, the approximate and the exact versions of MARS instantiations on CIFAR-100 dataset.
(a) AdamW-series
(b) Lion-series
(c) Shampoo-series
Figure 17: The test accuracy curves for the baselines of AdamW, Lion and Shampoo without variance reduction, the approximate and the exact versions of MARS instantiations on CIFAR-100 dataset.
D.5 Sensitivity to $\gamma$.
To explore the impact of $\gamma_{t}$, we test various $\gamma$s on GPT-2 small model, including constant and linearly changing schedules. And we plot the training and validation curves in Figure 18. It can be observed that there are slight differences among different $\gamma$s where 0.025 is the best $\gamma$. Therefore, we used $\gamma=0.025$ for other experiments in this paper.
Figure 18: Training loss and validation loss curves for MARS and MARS-approx on GPT-2 small (125M) models trained with MARS-AdamW-approx with different $\gamma$s, pretrained with OpenWebText dataset and plotted against training tokens.
Appendix E Hyper-parameter Settings
For training parameters, we did a grid search over learning rates between $\{1e-4,1.5e-4,3e-4,6e-4,1e-3,1.5e-3,3e-3,6e-3\}$, for weight decay coefficient, we did a grid search over $\{1e-1,1e-2,1e-3\}$. For AdamW baseline, although we utilized the golden standard learning rates in literature (a parameter search for AdamW have been done in Liu et al. (2023)), we also did a grid search on different parameters. Part of the results of different learning rates and learning rate schedules are also shown in Section 5.3.1 and 5.3.2, respectively. Table 10 summarizes the architectural hyperparameters for GPT-2 models with 125M (small), 355M (medium), 770M (large) and 1.5B (XL, only used in Appendix D.3) parameters. Table 11 lists the general hyperparameters used across all experiments, while Table 12 present the training hyperparameters for the small, medium, and large models, respectively.
@@ Appendix E Hyper-parameter Settings
Table 12: Hyper-parameters for GPT-2 experiments. We use $\gamma_{t}\equiv 0.025$ for MARS and $\mu=0.95$ for Muon.
Hyper-parameter GPT-2 Size AdamW Muon MARS-AdamW
Max learning rate small (125M) 6e-4 2e-2 6e-3
medium (355M) 3e-4 1e-2 3e-3
large (770M) 2e-4 6.67e-3 2e-3
Min learning rate small (125M) 3e-5 3e-5 3e-5
medium (355M) 6e-5 6e-5 6e-5
large (770M) 1e-5 1e-5 1e-5
$(\beta_{1},\beta_{2})$ small/medium/large (0.9,0.95) - (0.95,0.99)